OpenStudy (anonymous):

given the equation of the like y+10=-(x-15) what is the slope?

OpenStudy (anonymous):

First we need to put it in slope-intercept form. that is y=mx+b. So first distribute the - on the right side. Do you know how to do that?

OpenStudy (haseeb96):

y+10=-x+15 y=-x+15-10 y=-1x+5 here slope is -1

OpenStudy (mathmale):

Zeinas' approach will certainly work, and work fine. However, simply comparing the given equation to the point-slope form of the equation of a straight line would be faster: y - t0 = m (x - x0) y + 10 = -1(x - x0) It's immediately apparent that m = -1.

OpenStudy (anonymous):

That's a nice way to do it @mathmale. I didn't know how to do it like that. I learned something too lol

OpenStudy (anonymous):

How did you get 0 on the right side?

OpenStudy (anonymous):

Are you referring to the 0 in mathmale's reply?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

He used a simpler way. That is point-slope form. The form is y-y1=m(x-x1). And "m" is the slope. I'm not sure how he got the 0, but he just put the equation in that form. If it would have been y+10=-(x-15) The answer would have been the same. -1.

OpenStudy (anonymous):

I meant y+10=-1(x-15)

OpenStudy (mathmale):

Mathmale speaking: the standard point-slope form of the equation of a straight line could also be written as y-b = m (x-a), which represents the straight line through the point (a,b) with slope of m. Instead of (a,b), I used (x0, y0) to represent the given point. What this should read is "x sub 0, y sub 0," or, in other words, the "0" is nothing but a subscript.